# Extended OpenTT games dataset: A table tennis dataset for fine-grained shot type and point outcome detection

### v1

@@ 1 Introduction
A variety of machine learning approaches have been applied to table tennis video analysis. Foundational work such as TTNet[Voeikov_2020_CVPR_Workshops] focused on fundamental event spotting, using a multi-task encoder–decoder architecture to detect ball bounces and net hits from a sideline view, providing segmentation masks as a key output. Other research has shifted the focus to the players themselves. RallyTempPose[A_Stroke_of_Genius], for example, pioneered the prediction of future strokes by analyzing player skeletons, using a dual attention mechanism to capture game dynamics and opponent interactions. More directly related to stroke classification, researchers have explored various deep learning architectures. In the analogous domain of tennis, for example, a model such as SlowFast [feichtenhofer2019slowfastnetworksvideorecognition] exemplifies the effectiveness of temporal deep learning in general and is applied in [hovad2024classificationtennisactions] on the THETIS benchmark dataset [gourgari2013thetis].
For table tennis, ViSTec[vistec] represents a notable work, achieving high accuracy by leveraging a powerful VideoMae[VideoMAE] transformer model combined with a graph for contextual understanding. With regard to fine-grained stroke classification, this has been heavily driven by the introduction of specialized datasets. The most notable of these is TTStroke-21 [martin:hal-02937666], introduced as part of the MedieEval benchmarking initiative. With its annotations for 20 distinct stroke types, it now serves as the standard dataset for the associated benchmark test. In a different approach focusing on capturing high-fidelity motion, Kulkarni and Shenoy collected a custom dataset [kulkarni2021tabletennis]. Rather than using typical match footage, they recorded an isolated player from a head-on camera angle, arguing this provides a clearer view of stroke biomechanics for the 11 types they annotated. Some recent work has explored classifying strokes using only the ball’s trajectory data — a novel, minimalist approach that sidesteps visual analysis of the player entirely [kulkarni2023tabletennisstrokedetection]. While these works demonstrate significant value, a persistent challenge is data accessibility. The specialized datasets used in many of these studies, such as the one introduced by Kulkarni and Shenoy and the trajectory data for motion-based analysis, are not publicly released. Moreover, while benchmark datasets like TTStroke-21 are available for specific tasks, they often carry restrictive licenses that hinder broader reuse. This lack of a large-scale, openly licensed dataset with rich tactical labels limits reproducible research and comparative benchmarking.
@@ 2 Data
In this section, we describe the additions we made to the original OpenTTGames dataset[Voeikov_2020_CVPR_Workshops]. The code for accessing the extended data is provided in the link https://gitlab.compute.dtu.dk/emilh/table_tennis_data.
@@ 3.2 Fine-grained point outcome
In addition to the above annotations, we also decided to label how the rallies ended. Like for the stroke annotations, the rally outcomes are prefixed with the player who hit the rally-ending ball. The rally endings have been distributed into the following six categories, which should cover the various ways a rally may conclude, providing a detailed and fine-grained distinction between outcomes. Examples of the rally outcomes are shown in Figure 6.
  • (a) The out-class is defined relative to the left side. For the right side, the logic is mirrored.
@@
    Example: The left player hits the ball out
@@
  • (b) Net-class is defined relative to the left side. For the right side, the logic is mirrored.
@@
  • (c) winner-class is defined relative to the left side. For the right side, the logic is mirrored.
@@
  • (d) not_hitting_ball-class is defined relative to the left side. For the right side, the logic is mirrored.
@@
  • (e) double_bounce-class is defined relative to the right side. For the right side, the logic is mirrored.
@@
  • (f) miss_on_own_side-class is defined relative to the left side. For the right side, the logic is mirrored.
@@ 4 Conclusion
We presented an extended version of OpenTTGames with frame-accurate annotations for (i) side-aware shot classes based on the side of the table (left/right $\times$ {forehand, backhand}), (ii) player posture when the ball have contact with the bat (body lean and legs) and (iii) rally outcomes at point end Hereby we provide a table tennis stroke data set with detailed annotations but without an unclear license. By providing structured, frame-accurate stroke, posture, and outcome labels, our dataset enables supervision beyond event spotting and supports learning higher-level tactical patterns in table tennis play. All scripts for inspection and annotation are openly available, facilitating reproducibility and further extensions of the data.
Limitations and opportunities. The current release focuses on a single static viewpoint and match-level footage; posture labels are concise by design and do not yet capture full-body kinematics or racket-face orientation. Future extensions include (a) sound (the videos has no sound track) (b) new multi-view or mobile-camera recordings, (c) richer pose and contact attributes (e.g., contact height, stance width), (d) inter-rally context (serve patterns, receive strategies), and (e) stronger baselines that fuse ball, pose, and appearance streams. We hope this dataset and protocol catalyze open, comparable benchmarks for fine-grained shot type and point outcome modeling in racket sports.

### v2

@@ 1 Introduction
A variety of machine learning approaches have been applied to table tennis video analysis. Foundational work such as TTNet[10] focused on fundamental event spotting, using a multi-task encoder–decoder architecture to detect ball bounces and net hits from a sideline view, providing segmentation masks as a key output. Other research has shifted the focus to the players themselves. RallyTempPose[5], for example, pioneered the prediction of future strokes by analyzing player skeletons, using a dual attention mechanism to capture game dynamics and opponent interactions. More directly related to stroke classification, researchers have explored various deep learning architectures. In the analogous domain of tennis, for example, a model such as SlowFast [1] exemplifies the effectiveness of temporal deep learning in general and is applied in [4] on the THETIS benchmark dataset [2].
For table tennis, ViSTec[3] represents a notable work, achieving high accuracy by leveraging a powerful VideoMae[9] transformer model combined with a graph for contextual understanding. With regard to fine-grained stroke classification, this has been heavily driven by the introduction of specialized datasets. The most notable of these is TTStroke-21 [8], introduced as part of the MedieEval benchmarking initiative. With its annotations for 20 distinct stroke types, it now serves as the standard dataset for the associated benchmark test. In a different approach focusing on capturing high-fidelity motion, Kulkarni and Shenoy collected a custom dataset [7]. Rather than using typical match footage, they recorded an isolated player from a head-on camera angle, arguing this provides a clearer view of stroke biomechanics for the 11 types they annotated. Some recent work has explored classifying strokes using only the ball’s trajectory data — a novel, minimalist approach that sidesteps visual analysis of the player entirely [6]. While these works demonstrate significant value, a persistent challenge is data accessibility. The specialized datasets used in many of these studies, such as the one introduced by Kulkarni and Shenoy and the trajectory data for motion-based analysis, are not publicly released. Moreover, while benchmark datasets like TTStroke-21 are available for specific tasks, they often carry restrictive licenses that hinder broader reuse. This lack of a large-scale, openly licensed dataset with rich tactical labels limits reproducible research and comparative benchmarking.
@@ 2 Data
In this section, we describe the additions we made to the original OpenTTGames dataset[10]. The code for accessing the extended data is provided in the link https://github.com/moamal01/table_tennis_data.
@@ 3.2 Fine-grained point outcome
In addition to the above annotations, we also decided to label how the rallies ended. Like for the stroke annotations, the rally outcomes are prefixed with the player who hit the rally-ending ball. The rally endings have been distributed into the following six categories, which should cover the various ways a rally may conclude, providing a detailed and fine-grained distinction between outcomes. Examples of the rally outcomes are shown in Figure 6. The six categories are illustrated below for the left-side player (for the right-side player, the logic is mirrored).
  • (a) Out-class:
@@
    Example: The left player hits the ball out.
@@
  • (b) Net-class:
@@
  • (c) Winner-class
@@
  • (d) Not_hitting_ball-class:
@@
  • (e) Double_bounce-class:
@@
  • (f) Miss_on_own_side-class:
@@ 4 Conclusion
We presented an extended version of OpenTTGames with frame-accurate annotations of (i) left/right player and stroke type classes, (ii) player posture when the ball have contact with the bat (body lean and legs) and (iii) point outcome of the rally. Hereby we provide a table tennis stroke data set with detailed annotations and a transparent license. By providing structured, frame-accurate stroke, posture, and outcome labels, our dataset enables supervision beyond event spotting and supports learning higher-level tactical patterns in table tennis play. All scripts for inspection and annotation are openly available, facilitating reproducibility and further extensions of the data.